# REPLAY: Modeling Time-Varying Temporal Regularities of Human Mobility for Location Prediction over Sparse Trajectories

### v1

@@ Abstract
Location prediction forecasts a user’s location based on historical user mobility traces. To tackle the intrinsic sparsity issue of real-world user mobility traces, spatiotemporal contexts have been shown as significantly useful. Existing solutions mostly incorporate spatiotemporal distances between locations in mobility traces, either by feeding them as additional inputs to Recurrent Neural Networks (RNNs) or by using them to search for informative past hidden states for prediction. However, such distance-based methods fail to capture the time-varying temporal regularities of human mobility, where human mobility is often more regular in the morning than in other periods, for example; this suggests the usefulness of the actual timestamps besides the temporal distances. Against this background, we propose REPLAY, a general RNN architecture learning to capture the time-varying temporal regularities for location prediction. Specifically, REPLAY not only resorts to the spatiotemporal distances in sparse trajectories to search for the informative past hidden states, but also accommodates the time-varying temporal regularities by incorporating smoothed timestamp embeddings using Gaussian weighted averaging with timestamp-specific learnable bandwidths, which can flexibly adapt to the temporal regularities of different strengths across different timestamps. Our extensive evaluation compares REPLAY against a sizable collection of state-of-the-art techniques on two real-world datasets. Results show that REPLAY consistently and significantly outperforms state-of-the-art methods by 7.7%-10.9% in the location prediction task, and the bandwidths reveal interesting patterns of the time-varying temporal regularities.
@@ 1 Introduction
Location prediction is a key problem in human mobility modeling, serving as a fundamental building block for various location based services. The primary task is to predict a user’s future location based on the user’s historical mobility traces Noulas et al. (2012b). As a sequence modeling problem by nature, location prediction problems have been widely tackled by the literature using Recurrent Neural Networks (RNNs), such as vanilla RNN Mikolov et al. (2010), Long Short-Term Memory (LSTM) Hochreiter and Schmidhuber (1997) and Gated Recurrent Unit (GRU) Cho et al. (2014); these techniques have shown great success in language modeling to learn from word sequences. However, different from word sequences (i.e., text sentences), real-world user mobility traces are often sparse and incomplete, due to the nature of the data collection scheme and privacy issues. More precisely, on Location Based Social Networks (LBSNs) which are widely used as a primary data source for benchmarking location prediction techniques, users voluntarily share their spatiotemporal presence (check-ins) at a specific Point of Interest (POI) at a specific time; such a voluntary sharing protocol intrinsically implies sparse mobility traces. For example, an empirical study on a widely used Foursquare dataset shows that the median time between successive check-ins is about 16.72 hours Deng et al. (2023). Such sparsity and incompleteness hinder the application of RNNs to the location prediction problem.
@@
To tackle this sparsity issue, existing works strive to extend RNN architectures by incorporating spatiotemporal contexts, which have been shown as strong predictors for location prediction Gonzalez et al. (2008). A widely adopted approach is to add the spatiotemporal distances between successive check-ins as additional inputs together with sequences of POIs to RNN units Liu et al. (2016); Neil et al. (2016); Zhu et al. (2017); Kong and Wu (2018); Cui et al. (2019); Zhao et al. (2019, 2020). Despite its wide adoption, recent studies have shown that this approach cannot fully capture the temporal periodicity and spatial regularity of human mobility patterns Yang et al. (2020a), as considering spatiotemporal distances between check-ins that are several steps apart (i.e., high-order distances) in the check-in sequences are often more informative than those between successive check-ins for location prediction. Subsequently, a flashback mechanism has been introduced to explicitly use such high-order spatiotemporal distances to search past hidden states with high predictive power (i.e., historical hidden states with similar spatiotemporal contexts to the current one) Yang et al. (2020a); Li et al. (2021a); Cao et al. (2021); Rao et al. (2022); Liu et al. (2022); Wu et al. (2022); Ye et al. (2023). However, despite the improved performance, the temporal distances still fail to capture the time-varying temporal regularities of human mobility, as we discuss below.
@@
Temporal regularity has been revealed as a universal law of human mobility Gonzalez et al. (2008); Noulas et al. (2012a), which can be evidenced by the periodicity of human activities even in sparse human mobility traces Cho et al. (2011). Figure 1 illustrates the returning probability of user check-ins over time on the Foursquare dataset, which is defined as the probability of a user re-checking in at a POI a certain period of time after her first check-in at that POI Gonzalez et al. (2008); Lian et al. (2015). We observe an obvious daily (periodic) revisiting pattern, which implies that the historical check-ins with a temporal distance closer to these daily peaks have higher predictive power; this serves as the foundation for the flashback mechanism Yang et al. (2020a). In this paper, we further reveal that the strength of the temporal regularities varies across different time periods during a day (or across different days in a week, e.g., weekdays v.s. weekends). On the Foursquare dataset, the mobility pattern in the daytime is mostly attributed to working-related activities, which is usually more regular than the nighttime mobility pattern that is mostly attributed to entertainment-related activities [^1] Note that user check-ins on LBSNs are mostly for social sharing purposes, where check-ins at entertainment-related POIs are much more than those at home-related POIs during the nighttime Yang et al. (2020b).. Figure 1 also compares the daytime and nighttime returning probabilities, where the daytime returning probability is significantly higher than nighttime, implying that the temporal regularity is much stronger in the daytime than in the nighttime. This implies that the hidden state 24 hours back is more informative for location prediction if the current time is daytime rather than nighttime. However, such time-varying temporal regularities cannot be captured by the distance-based methods (i.e., merely considering the temporal distances), and thus require involving the actual timestamps of check-ins.
@@
Motivated by this observation, we propose REPLAY, a general RNN based architecture learning to capture the vaRying tEmPoral reguLAritY of human mobility for location prediction. Specifically, REPLAY is designed to seamlessly incorporate smoothed timestamp embeddings with learnable bandwidths into the flashback mechanism. More precisely, for each check-in in a check-in sequence, we first transform its time to an hour-in-week timestamp $t_{n}$, where $1\leq n\leq 168$ (one of the 168 hours in a week) and then generate a smoothed timestamp embedding using Gaussian weighted averaging with a timestamp-specific learnable bandwidth $\sigma_{n}$ (controlling the width of the Gaussian weighting function centered at the timestamp $n$). Afterward, we concatenated the smoothed timestamp embedding with the corresponding POI embedding, which is then fed into an RNN with the flashback mechanism, where the spatiotemporal distances in sparse trajectories are used to search for the informative past hidden states. Finally, we combine the hidden state output by the RNN, a user embedding, and a smoothed timestamp embedding generated from a query time to make predictions on future POIs. The timestamp-specific learnable bandwidths here can automatically learn to adapt to the temporal regularities of different strengths across different timestamps. Intuitively, a smaller $\sigma_{n}$ leads to a peaky Gaussian, which implies a stronger regularity at timestamp $n$, as the smoothed timestamp embedding relies mostly on the embedding of $t_{n}$ itself and less on the information from its neighboring timestamps for prediction. This is also evidenced by our experiments later; the learnt bandwidths for the timestamps in the daytime are indeed 7.5% smaller on average than those in the nighttime, which well corresponds to our observations in Figure 1. Our contributions are summarized as follows:
@@
  • We conduct a thorough evaluation using two real-world LBSN datasets. Results show that our REPLAY outperforms a sizable collection of state-of-the-art location prediction methods, with 7.7%-10.9% improvement over the best-performing baselines. Moreover, the learnt bandwidths reveal interesting temporal regularity patterns: 1) morning mobility shows a stronger regularity compared to other time periods, which is consistent on both weekdays and weekends; 2) weekend mobility is less regular than weekdays in general, while the weekend afternoon sometimes shows the weakest regularity.
@@ 2 Related Work
Location prediction is one key task of mobility modeling, which predicts a user’s future location based on her historical mobility traces. Traditional methods mainly focus on various mobility features including historical visit counts Noulas et al. (2012b); Gao et al. (2012), activity preferences Ye et al. (2013); Yang et al. (2015), social influences Cho et al. (2011); Sadilek et al. (2012), or automatically-learnt features (embedding) Xie et al. (2016); Feng et al. (2017); Qian et al. (2019); Yang et al. (2019). However, these techniques have limited capability of capturing the sequential patterns of user mobility, which have been shown as an important predictor for location prediction Liu et al. (2016).
To model the sequential patterns of user mobility, early methods often leverage Markov Chains to estimate a transition matrix encoding the probability of visiting a POI conditioned on past visits. For example, for location prediction/recommendation problems, transition matrices are built with factorized personalized Markov Chains with spatial constraints Cheng et al. (2013); Feng et al. (2015); Zhang et al. (2014a). However, these methods have intrinsic limitations in capturing the long-term dependencies of visited locations in user mobility traces.
@@
To capture the long-term dependencies, deep learning sequential models (e.g., RNN, LSTM, GRU, and Attention Neural Networks) have been widely used in location prediction tasks. To handle real-world mobility traces that are often sparse and incomplete, spatiotemporal contexts are often considered. A widely adopted approach is to add the spatiotemporal distances between successive check-ins as additional inputs to the RNN units. For example, the spatial distances between successively checked POIs are fed as additional inputs to RNNs Cui et al. (2019); STRNN Liu et al. (2016) designs RNN transition matrices parameterized by the spatiotemporal distances between successive check-ins; HST-LSTM Kong and Wu (2018) extends existing gates in LSTMs to allow the spatiotemporal distance as an additional input; STGN Zhao et al. (2019) incorporates additional gates controlled by the spatiotemporal distances to LSTMs; NeuNext Zhao et al. (2020) extend STGN via POI context prediction to help next POI recommendation tasks. STAN Luo et al. (2021) uses spatiotemporal attention networks to capture the dependencies between non-adjacent check-ins; GeoSAN Lian et al. (2020) uses hierarchical gridding of GPS locations for spatial discretization and uses self-attention layers for matching. Despite the improved performance of these methods, recent studies have shown that considering the spatiotemporal distances between successive check-ins cannot fully capture the temporal periodicity and spatial regularity of human mobility patterns Yang et al. (2020a), and proposed a flashback mechanism to explicitly use the high-order spatiotemporal distances to search past hidden states with high predictive power (i.e., historical hidden states that share similar spatiotemporal contexts to the current one). For example, Flashback Yang et al. (2020a) aggregates past hidden states with high predictive power using a context-aware weighting mechanism; BiGRU Cao et al. (2021) designs a bi-directional GRU with the flashback mechanism; BSDA Li et al. (2021a) designs a bi-direction speculation method for location prediction, combining flashback with an additional RNN which models POIs’ appeal to users; Bi-STAN Wu et al. (2022) extends flashback mechanism for missing check-in imputation; RTPM Liu et al. (2022) combines flashback with real-time user preference models for real-time location recommendation. In addition, on top of the sequence models, researchers also resort to graph-based models to help the location prediction performance, where the graph-based models can easily integrate additional information beyond trajectory data (e.g., category or user social relationship), such as graph-enhanced attention networks Li et al. (2021b); Wang et al. (2022b, a), graph-enhanced flashback networks Rao et al. (2022), and hierarchical multi-task graph recurrent networks Lim et al. (2022). However, all of these existing methods overlook the time-varying temporal regularities of user mobility (as evidenced in Figure 1), which motivates us to design REPLAY as presented below.
@@ 3 REPLAY
To capture the time-varying temporal regularities of human mobility, we design REPLAY to seamlessly incorporate smoothed timestamp embeddings with learnable bandwidths into the flashback mechanism. Figure 2 shows the overview of our REPLAY. Given a user’s mobility trace denoted as a POI sequence $\{...,p_{i-2},p_{i-1},p_{i},...\}$ and its corresponding check-in time $\{...,t_{i-2},t_{i-1},t_{i},...\}$, where $i$ denote the order of check-in in the mobility trace, our objective is to predict a future POI $p_{i+1}$ that a user will probably visit at a query time $t_{i+1}$. To this end, REPLAY first generates a smoothed timestamp embedding for each check-in time in the sequence, and then concatenates it with the corresponding POI embedding which are then fed into an RNN with the flashback mechanism. Finally, we combine the hidden state output by the RNN, a user embedding, and a smoothed timestamp embedding generated from the query time to make predictions on future POIs. We present each component in detail below.

### v3

@@ Abstract
Next-location prediction aims to forecast which location a user is most likely to visit given the user’s historical data. As a sequence modeling problem by nature, it has been widely addressed using Recurrent Neural Networks (RNNs). To tackle the intrinsic sparsity issue of real-world user mobility traces, spatiotemporal contexts have been shown as significantly useful. Existing solutions mostly incorporate spatiotemporal distances between locations in mobility traces, either by integrating them into the RNN units as additional information, or utilizing them to search for informative historical hidden states to improve prediction. However, such distance-based methods fail to capture the time-varying temporal regularities of human mobility, where human mobility is often more regular in the morning than in other time periods, for example; this suggests the usefulness of the actual timestamps besides the temporal distances. Under this circumstance, we propose REPLAY, learning to capture the time-varying temporal regularities for location prediction based on general RNN architecture. Specifically, REPLAY is designed on top of a flashback mechanism, where the spatiotemporal distances in sparse trajectories are used to search for the informative past hidden states; to accommodate the time-varying temporal regularities, REPLAY incorporates smoothed timestamp embeddings using Gaussian weighted averaging with timestamp-specific learnable bandwidths, which can flexibly adapt to the temporal regularities of different strengths across different timestamps. We conduct a comprehensive evaluation, comparing REPLAY against a wide range of state-of-the-art methods. Experimental results show REPLAY significantly and consistently outperforms state-of-the-art methods by 7.7%-10.9% in the location prediction task, and the learnt bandwidths reveal interesting patterns of the time-varying temporal regularities.
Index Terms:
User mobility, Sparse trajectory, Location prediction, Recurrent neural networks
@@ 1 Introduction
Location prediction plays a crucial role in the modeling of human mobility, serving as a fundamental building block for various location based services. The primary aim is to forecast which location a user is most likely to visit given the user’s historical data [1]. As a sequence modeling problem by nature, location prediction problems have been widely tackled by the literature using Recurrent Neural Networks (RNNs), such as vanilla RNN [2], Gated Recurrent Unit (GRU) [3] and Long Short-Term Memory (LSTM) [4]; these techniques have shown great success in modeling language data. However, different from word sequences (i.e., text sentences), real-world user mobility traces are often sparse and incomplete, due to the nature of the data collection scheme and privacy issues. More precisely, Location Based Social Networks (LBSNs) are valuable data sources for benchmarking location prediction techniques, where users voluntarily share their spatiotemporal presence (check-ins) at specific Points of Interest (POIs) and specific times. However, due to the voluntary sharing, the resulting mobility traces tend to be sparse in nature. For example, an empirical investigation using the commonly employed Foursquare dataset shows that the median time interval between consecutive check-ins is approximately 16.72 hours [5]. This sparsity and incompleteness in data present challenges for applying RNNs to the task of location prediction.
@@
To tackle this sparsity issue, existing works strive to extend RNN architectures by incorporating spatiotemporal contexts, which have been shown as strong predictors for location prediction [6]. A widely adopted approach is to add the spatiotemporal distances between successive check-ins as additional inputs together with sequences of POIs to RNN units [7, 8, 9, 10, 11, 12, 13]. Despite its wide adoption, recent studies have shown that this approach cannot fully capture the temporal periodicity and spatial regularity of human mobility patterns [14], as considering spatiotemporal distances between check-ins that are several steps apart (i.e., high-order distances) in the check-in sequences are often more informative than those between successive check-ins for location prediction. Subsequently, a flashback mechanism has been introduced to explicitly use such spatiotemporal distances in sparse trajectories search for the informative past hidden states (i.e., historical hidden states with similar spatiotemporal contexts to the current one) [14, 15, 16, 17, 18, 19, 20, 21, 20]. However, despite the improved performance, the temporal distances still fail to capture the time-varying temporal regularities of human mobility, as we discuss below.
@@
Temporal regularity has been revealed as a universal law of human mobility [6, 22], which can be evidenced by the periodicity of human activities even in sparse human mobility traces [23]. Figure 1 shows the temporal returning probability of user check-ins on the Foursquare dataset. This probability represents the probability of users revisiting a particular POI within a specific time interval after their initial check-in at that POI [6, 24]. We observe an obvious daily (periodic) revisiting pattern, which implies that historical check-ins closer to these daily peaks in temporal distance have higher predictive power; this serves as the foundation for the flashback mechanism [14]. In this paper, we further reveal that the strength of the temporal regularities varies across different time periods during a day (or across different days in a week, e.g., weekdays v.s. weekends). On the Foursquare dataset, the mobility pattern in the daytime is mostly attributed to working-related activities, which is usually more regular than the nighttime mobility pattern that is mostly attributed to entertainment-related activities [^1] Note that user check-ins on LBSNs are mostly for social sharing purposes, where check-ins at entertainment-related POIs are much more than those at home-related POIs during the nighttime [25].. Figure 1 also compares the daytime and nighttime returning probabilities, where the daytime returning probability is significantly higher than nighttime, implying that the temporal regularity is much stronger in the daytime than in the nighttime. This implies that the hidden state 24 hours back is more informative for location prediction if the current time is daytime rather than nighttime. However, such time-varying temporal regularities cannot be captured by the distance-based methods (i.e., merely considering the temporal distances), and thus require involving the actual timestamps of check-ins.
@@
Motivated by this observation, we propose REPLAY, a general RNN based architecture learning to capture the vaRying tEmPoral reguLAritY of human mobility for location prediction. Specifically, REPLAY is designed to seamlessly incorporate smoothed timestamp embeddings with learnable bandwidths into the flashback mechanism. More precisely, for each check-in in a check-in sequence, we first transform its time to an hour-in-week timestamp $t_{n}$, where $1\leq n\leq 168$ (one of the 168 hours in a week) and then generate a smoothed timestamp embedding using Gaussian weighted averaging with a timestamp-specific learnable bandwidth $\sigma_{n}$ (controlling the width of the Gaussian weighting function centered at the timestamp $n$). Afterward, we concatenated the smoothed timestamp embedding with the corresponding POI embedding, which is then fed into an RNN with the flashback mechanism, where the spatiotemporal distances in sparse trajectories are used to search for the informative past hidden states. Finally, we combine the hidden state output by the RNN, a user embedding, and a smoothed timestamp embedding generated from a query time to make predictions on future POIs. The timestamp-specific learnable bandwidths here can automatically learn to adapt to the temporal regularities of different strengths across different timestamps. Intuitively, a smaller $\sigma_{n}$ leads to a peaky Gaussian, which implies a stronger regularity at timestamp $n$, as the smoothed timestamp embedding relies mostly on the embedding of $t_{n}$ itself and less on the information from its neighboring timestamps for prediction. This is also evidenced by our experiments later; the learnt bandwidths for the timestamps in the daytime are indeed 7.5% smaller on average than those in the nighttime, which well corresponds to our observations in Figure 1. We summarize our contributions as follows:
@@
  • We conduct a thorough evaluation using two real-world LBSN datasets. Results show that our REPLAY outperforms a sizable collection of cutting-edge location prediction methods, with 7.7%-10.9% improvement over the best-performing baselines. Moreover, the learnt bandwidths reveal interesting temporal regularity patterns: 1) morning mobility shows a stronger regularity compared to other time periods, which is consistent on both weekdays and weekends; 2) weekend mobility is less regular than weekdays in general, while the weekend afternoon sometimes shows the weakest regularity.
@@ 2 Related Work
2.1 Human Mobility Trajectory
Ravenstein published work on the analysis of migration patterns using human mobility trajectories derived from demographic data [26] is pioneering work in studying human mobility. Nowadays, the widespread use of smart devices equipped with sensors has significantly enhanced the accessibility of human trajectory data for comprehensive studies [27]. Human trajectory data can be divided into two different types based on the data collection method as follows.
Firstly, there are continuously sampled trajectories, which encompass regularly recorded location sequences obtained from devices equipped by individuals. For instance, Microsoft Research Asia conducted the GeoLife experiment [28] which collects trajectories from 182 volunteers for more than three years by capturing GPS traces every five seconds or ten meters. Another notable example is that IDIAP Research Institute and Nokia Research Center conducted the Lausanne Data Collection Campaign [29], where the mobility traces of 200 individuals were collected. Each individual carried a smartphone that periodically recorded various sensor readings (e.g., Bluetooth, WLAN) at intervals such as every 60 seconds. While these human trajectory datasets provide detailed information about individual mobility patterns, their scale is often limited due to designed settings and individuals’ concerns regarding privacy (e.g., unwillingness to install software that continuously monitors and collects data on personal devices).
Secondly, there are voluntarily shared trajectories, which consist of self-reported location sequences, primarily provided by users on social media network platforms. For instance, a large number of individuals willingly share their check-ins on LBSNs, contributing to the accumulation of a substantial collection of human trajectories. However, people often choose to disclose a portion of their check-ins by reporting at specific POIs. This means that users may visit various POIs without recording, influenced by factors such as their preferences for certain places, simply forgetting to do so, or even privacy concerns. Consequently, despite LBSNs being commonly acknowledged as an important data source for conducting extensive research and data collection, the resulting mobility trajectories inherently exhibit sparsity [30, 31, 32]. This sparsity needs to be taken into careful consideration when conducting data analysis tasks [33]. In light of this context, the focus of this paper is to investigate the problem of location prediction over such sparsely populated trajectories.
2.2 Location Prediction
The spatiotemporal regularity of human mobility has been revealed with ample empirical evidence in the literature [6]. Universal mobility patterns have been discovered across different cities using statistical methods [22, 34], which serve as the foundation for human mobility modeling. Location prediction plays an important role in human mobility studying, which aims to predict the next location of users that they are most likely to visit by leveraging their historical traces. Traditional methods mainly focus on various mobility features including activity preferences [35, 36], historical visit counts [1, 37], social information [23, 38], or features (embedding) learnt by graph embedding techniques [39, 40, 41, 32]. In addition, factorization and generative methods have been utilized to tackle the challenges of location recommendation and prediction tasks [42, 43, 44, 45, 46]. However, these techniques possess limited ability to capture the sequential patterns that are inherent in human mobility. , which have been proven to be crucial predictors for accurate location prediction [7].
To model the sequential patterns of human mobility, (Hidden) Markov Chain based methods have been extensively employed for sequential prediction [47, 48, 49]. The fundamental idea is to calculate the transition probability of a particular state according to the preceding one. Factorizing Personalized Markov Chains (FPMC) is a typical technique employed in this context [50], which factorizes a personalized transition matrix in user-specific latent factors. To address the location prediction problem, FPMC has been extended by integrating spatial constraints into its modeling framework [48, 49]. However, one limitation of these location prediction techniques is their inability to effectively capture long-term dependencies in human mobility trajectories.
@@ 2.2 Location Prediction
To capture the long-term dependencies, Recurrent Neural Networks (e.g., RNN, GRU and LSTM) have been widely used in location prediction tasks. To handle real-world mobility traces that are often sparse and incomplete, spatiotemporal contexts are often considered in the RNNs. A common approach widely adopted is to augment the Recurrent Neural Network (RNN) units with the inclusion of temporal and spatial distances between consecutive check-ins as extra information. For instance, Distance2Pre [11] considers the spatial distance between consecutive check-ins as extra information; STRNN [7] extends RNN with time and spatial-specific transition matrices for location prediction; HST-LSTM [10] extends LSTMs gates to consider spatiotemporal distance; STGN [12] adds spatial and temporal gates to the conventional LSTM to capture sequential patterns of users; NeuNext [13] jointly learn the location context prediction and the next location prediction tasks. STAN [51] uses spatiotemporal attention networks to capture the dependencies between non-adjacent check-ins; GeoSAN [52] uses hierarchical gridding of GPS locations for spatial discretization and uses self-attention layers for matching. Despite the improved performance of these methods, we have shown in our previous study [14] that only considering the spatiotemporal distances between consecutive check-ins cannot fully capture the two universal human mobility laws, i.e., spatial regularity and temporal periodicity, for location prediction.
In this context, the flashback mechanism proposed by our previous work [14] explicitly utilizes high-order spatiotemporal distance information to find informative past hidden states. This flashback mechanism has been widely utilized by subsequent studies. For instance, BiGRU [16] incorporates the flashback network with a bi-directional GRU; BSDA [15] incorporates the flashback mechanism with an extra RNN which models correlations of user appealed by POIs from bi-direction speculation; Bi-STAN [19] utilizes flashback mechanism for missing value imputation task; RTPM [18] addresses real-time location recommendation task by combining flashback mechanism with real-time user preference; Graph-Flashback [17] incorporates knowledge graph embeddings technique into the flashback framework. Note that apart from RNNs, other sequential methods have also been utilized for location prediction, such as geography-aware self-attention networks [52], graph-enhanced attention networks [53, 54], spatiotemporal attention networks [51], and hierarchical multi-task graph recurrent networks [55]. However, all of these existing methods overlook the time-varying temporal regularities of user mobility (as evidenced in Figure 1), which motivates us to design REPLAY that can automatically adapt to the temporal regularities of different strengths over time.
Compared to our previous work Flashback [14], the current work REPLAY makes the following differences. First, beyond the temporal periodicity, we further reveal the time-varying temporal regularities of human mobility in this work. Second, to capture the time-varying temporal regularities for location prediction, we design REPLAY on top of Flashback by seamlessly incorporating smoothed timestamp embeddings with learnable bandwidths into RNNs with the Flashback mechanism, where Flashback is now one of the three building blocks of REPLAY (more detail below). Third, the new experiments show superior performance of REPLAY against state-of-the-art techniques; in particular, compared to our previous work Flashback, REPLAY shows a significant improvement of 21.8%-43.9% in location prediction tasks.
@@ 3 REPLAY
Motivated by the observation of the time-varying temporal regularities of human mobility, we design REPLAY to seamlessly incorporate smoothed timestamp embeddings with learnable bandwidths into the flashback mechanism. The timestamp-specific learnable bandwidths can automatically adapt to the temporal regularities of different strengths across different timestamps. Figure 2 shows the overall architecture of our proposed REPLAY. Given an individual’s trace denoted as a sequence of POIs $\{...,p_{i-2},p_{i-1},p_{i},...\}$ and its associated check-in time $\{...,t_{i-2},t_{i-1},t_{i},...\}$, where $i$ denote the order of check-in in the mobility trace, our objective is to predict a future POI $p_{i+1}$ that an individual will probably visit at a query time $t_{i+1}$. To this end, REPLAY first generates a smoothed timestamp embedding for each check-in time in the sequence, and then concatenates it with the corresponding POI embedding which is then fed into an RNN with the flashback mechanism. Finally, we combine the hidden state output by the RNN, a user embedding, and a smoothed timestamp embedding generated from the query time to make predictions on future POIs conditioned on the query time. We present each component in detail below.